OpenStudy (anonymous):

The arithmetic mean of the ages of 20 boys is 13 years, 4 months and 5 days. Find the sum of their ages. One of the boys is of exactly 15 years. What is the average age of the remaining boys?

OpenStudy (mertsj):

Change 13 years, 4 months and 5 days to years. Multiply that answer by 20 and that will be the sum of their ages. Now subtract 15 and divide the result by 19, and that will be the average age of the remaining boys.

OpenStudy (mertsj):

13 years + 1/3 year + 1/73 year = 13 76/219 years. 20(13 76/219) = 261 121/219 years

OpenStudy (mertsj):

261 121/219 - 15 = 246 121/219

OpenStudy (anonymous):

I would like to ask how did you change 5 days into years? diveide by 36 right? 5/360 ? = 1/72

OpenStudy (anonymous):

divide*

OpenStudy (mertsj):

246 121/219 divided by 19 = 12 4063/4161

OpenStudy (mertsj):

5/365

OpenStudy (anonymous):

Why 365?

OpenStudy (mertsj):

There are 365 days in 1 year

OpenStudy (anonymous):

Oh ohkay.

OpenStudy (mertsj):

Before computers, in the business world, they used 1 year = 360 days but not any more.

OpenStudy (anonymous):

I'm getting 13.26 as the answer.

OpenStudy (anonymous):

Thank you for the help Mertsj and Happy new year :)

OpenStudy (mertsj):

I got 12 .98 but I did not check my calculations yet.

OpenStudy (mertsj):

Also, do you have to give the answer in years, months, and days as it was given you?

OpenStudy (anonymous):

No, it's just the average. But if I want to express in months years and days how do i?

OpenStudy (mertsj):

Yep. I redid it and got 13.26 just as you did.

OpenStudy (anonymous):

Okay now if I want to express in months years and days how do i?

OpenStudy (mertsj):

Take .26 times 365 and that would be the number of days. Then change that to months and days.

OpenStudy (anonymous):

13 years 3 months and 5 days.

OpenStudy (mertsj):

Good job!!

OpenStudy (anonymous):

Thanks. :)

OpenStudy (mertsj):

Happy New Year.

OpenStudy (anonymous):

Same to you :)

OpenStudy (anonymous):

happy new year